LASER ABSORPTION AND EMISSION SPECTROSCOPY OF THE YELLOW GREEN BAND SYSTEM OF TiCl^{+}

\underline{LEONID A. KALEDIN} AND MICHAEL C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

The Yellow-Green band system of TiCl^{+} has been studied using a dual-beam laser absorption technique in a hollow cathode discharge^{a}, and conventional emission spectroscopy. The quartz discharge cell for emission spectroscopy consisted of a 5 cm long, 2 mm internal diameter capillary similar to that used for a velocity modulation emission spectroscopy^{b}. The discharge was modulated at 25-125 kHz by a power supply which produced a bipolar sine wave. Samples of TiCl_{4} in Ar were flowed slowly through the cell from the center of the capillary and symmetrically pumped out via the electrodes. The Yellow-Green band system was observed previously in the emission^{c} and absorption^{d} spectra of TiCl^{+}. Molecular parameters and internuclear distances for the [17.9]^{3}Δ and X^{3}Φ states of the ^{48}Ti^{35}Cl^{+} and ^{48}Ti^{37}Cl^{+} isotopomers will be presented.

Work supported by AFOSR and the Mission Research Corporation.

aL. A. Kaledin, A. L. Kaledin, and M. C. Heaven, J. Mol. Spectrosc. 179, 246-252 (1996).

bW. Y. Fan and P. A. Hamilton, Chem. Phys. Lett. 230, 555-560 (1994).

cW. J. Balfour and K. S. Chandrasekhar, J. Mol. Spectrosc. 139, 245-252 (1990).

dL. A. Kaledin, J. P. Parrish, and M. C. Heaven, 51-st International Symposium on Molecular Spectroscopy, Paper ME08 (1996)